find N base 0 and k in the exponential function N(t)= N base 0e^kt, given that N(0) = 11 and N(4) = 5. state k acurate to six decimal places
Answer by stanbon(75887)

find No and k in the exponential function N(t)= No e^kt, given that N(0) = 11 and N(4) = 5. state k accurate to six decimal places
------------------
N(t)= No e^kt
--------------
N(0)=No e^(k*0)
N(0)=No e^0
N(0)=No = 11
---------------
N(4)=11 e^(4k)
5=11 e^(4k)
5/11 = e^(4k)
Take the natural log of both sides to get:
4k= -0.7884573604...
k= -0.1971143401...
-----------
Cheers,
Stan H.